shasha12:

Given that the surface area of a cube is given by the formula SA=6s2, what is the value of SA if s=1/2 cm? A.1 cm2 B.3 cm2 C.1/2 cm2 D.1 1/2 cm2

shasha12:

I think the answer is A

563blackghost:

So you have the equation: \(\bf{SA=6(\frac{1}{2})^{2}}\) First you would apply the exponent. Since it is a fraction you raise the denominator by the exponent. \(\bf\large{\frac{1}{2^{2}} = \frac{1}{4}}\) In which you multiply by 6. \(\large\bf\color{red}{6 \times \frac{1}{4}}\)

shasha12:

so the answer is 1 1/2

563blackghost:

Correct.

shasha12:

thankings you

563blackghost:

np